# Kinematics of Circumgalactic OĀ VIĀ Gas and Disk Rotation of Star-forming Galaxies

### v1

@@ 5 Conclusion
We presented a compilation of new and archival HST/COS observations of quasars behind , star-forming galaxies. All 18 quasar sightlines probe the CGM within 45āĀ of the major axes of the host galaxies with inclined disks (ā), doubling the existing number of major-axis sightlines in the literature for studying the relationship between the OĀ VIĀ and disk rotation kinematics.
@@
We compared the kinematics of the OĀ VIĀ gas with that of the low-ionization-state gas (traced by SiĀ IIĀ and SiĀ III) and disk rotation. We showed that while individual velocity OĀ VIĀ components do not correlate with disk rotation, the bulk of OĀ VIĀ gas, as measured by the column-density-weighted OĀ VIĀ velocity in each sightline, rarely counter-rotates with the galaxy disk (FigureĀ 5). In general, SiĀ IIĀ and SiĀ IIIĀ gas are less likely to counter-rotate with the disk compared to OĀ VIĀ (FiguresĀ 6 and 7). But more importantly, we matched velocity components between OĀ VIĀ and low ions (SiĀ IIĀ and SiĀ III) by minimizing the difference of their velocity centroids. We showed that the OĀ VIĀ velocity components with matches with low ions were mainly found in sightlines at small impact parameters of below 100 kpc, whereas those OĀ VIĀ velocity components without low-ion matches were on average at larger impact parameters. These low-ion-matched OĀ VIĀ velocity components are also more frequently corotate with the disk compared to those OĀ VIĀ velocity components without a low-ion match (FigureĀ 9).
@@
Previous literature has shown that low-ionization-state gas is commonly detected in the inner CGM, and major-axis sightlines often detect such gas to corotate with the galaxy disk. Although absorption-line measurements do not reveal where the gas lies along the line of sight, our results of the low-ion matched OĀ VIĀ velocity components sharing these properties suggest that these OĀ VIĀ components trace the gas that is co-spatial with that of the low ions, i.e., on the extended disk plane in the inner CGM. In contrast, the OĀ VIĀ velocity components without a low-ion match traces gas at large 3D radius. Furthermore, if some of the OĀ VIĀ gas resides on the extended disk plane, we asked the question that whether the gas has sufficient angular momentum to stay on circular orbits. If not, this could suggest the gas is infalling. We showed that about 45% of the OĀ VIĀ velocity components lack sufficient angular momentum to maintain a circular orbit. This fraction increases with decreasing ionization potential, with 80% of SiĀ IIĀ velocity components tracing radially infalling gas (FigureĀ 11). Interestingly, repeating this calculation using the column-density-weighted velocities increases the inflow fractions of OĀ VIĀ and SiĀ IIIĀ also to 80%. We interpret this result as a consequence of the turbulent motion adding random angular momentum vectors to the gas clouds associated with individual velocity components. Therefore, the corotation and inflow signatures are more likely to be observationally detected only when we consider the average angular momentum of the gas using the column-density-weighted velocities. Our future work will include photoionization modeling and geometric models for explaining the circumgalactic gas flows of OĀ VIĀ and low-ionization-state gas and to understand the CGM angular momentum and gas accretion onto galaxies.
@@
This work was supported by NASA through HST-GO-15866.002-A from the Space Telescope Science Institute. We gratefully acknowledge the support for SHH from the National Science Foundation under AST-2202328. JS was supported by the Israel Science Foundation (grant No.Ā 2584/21). Some observations reported here were obtained at the MMT Observatory, a joint facility of the University of Arizona and the Smithsonian Institution. Some of the data presented herein were obtained at the W.Ā M.Ā Keck Observatory, which is operated as a scientific partnership among the California Institute of Technology, the University of California and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. Some observations were supported by Swinburne Keck programmes 2014B W018E, 2015 W187E, and 2016A W032E, W056E. Some presented data are based on observations obtained with the Apache Point Observatory 3.5 m telescope, which is owned and operated by the Astrophysical Research Consortium.
@@ Appendix A Absorption System Associated with Individual Galaxies
This Appendix includes the Voigt profile fits of absorption associated with the target galaxies in individual sightlines. FigureĀ 13 shows the Voigt profile fits, and TableĀ 6 lists the parameters of the fits. Because not all ions were used in the analysis of this paper, we only present the measurements in the Appendix.

### v2

@@ 5 Conclusion
We presented a compilation of new and archival HST/COS observations of quasars behind , star-forming galaxies. All 18 quasar sightlines probe the CGM within 45ā of the major axes of the host galaxies with inclined disks (ā), doubling the existing number of major-axis sightlines in the literature for studying the relationship between the OĀ VI and disk rotation kinematics.
@@
We compared the kinematics of the OĀ VI gas with that of the low-ionization-state gas (traced by SiĀ II and SiĀ III) and disk rotation. We showed that while individual velocity OĀ VI components do not correlate with disk rotation, the bulk of OĀ VI gas, as measured by the column-density-weighted OĀ VI velocity in each sightline, rarely counter-rotates with the galaxy disk (FigureĀ 5). In general, SiĀ II and SiĀ III gas are less likely to counter-rotate with the disk compared to OĀ VI (FiguresĀ 6 and 7). But more importantly, we matched velocity components between OĀ VI and low ions (SiĀ II and SiĀ III) by minimizing the difference of their velocity centroids. We showed that the OĀ VI velocity components with matches with low ions were mainly found in sightlines at small impact parameters of below 100 kpc, whereas those OĀ VI velocity components without low-ion matches were on average at larger impact parameters. These low-ion-matched OĀ VI velocity components are also more frequently corotate with the disk compared to those OĀ VI velocity components without a low-ion match (FigureĀ 9).
@@
Previous literature has shown that low-ionization-state gas is commonly detected in the inner CGM, and major-axis sightlines often detect such gas to corotate with the galaxy disk. Although absorption-line measurements do not reveal where the gas lies along the line of sight, our results of the low-ion matched OĀ VI velocity components sharing these properties suggest that these OĀ VI components trace the gas that is co-spatial with that of the low ions, i.e., on the extended disk plane in the inner CGM. In contrast, the OĀ VI velocity components without a low-ion match traces gas at large 3D radius. Furthermore, if some of the OĀ VI gas resides on the extended disk plane, we asked the question that whether the gas has sufficient angular momentum to stay on circular orbits. If not, this could suggest the gas is infalling. We showed that about 45% of the OĀ VI velocity components lack sufficient angular momentum to maintain a circular orbit. This fraction increases with decreasing ionization potential, with 80% of SiĀ II velocity components tracing radially infalling gas (FigureĀ 11). Interestingly, repeating this calculation using the column-density-weighted velocities increases the inflow fractions of OĀ VI and SiĀ III also to 80%. We interpret this result as a consequence of the turbulent motion adding random angular momentum vectors to the gas clouds associated with individual velocity components. Therefore, the corotation and inflow signatures are more likely to be observationally detected only when we consider the average angular momentum of the gas using the column-density-weighted velocities. Our future work will include photoionization modeling and geometric models for explaining the circumgalactic gas flows of OĀ VI and low-ionization-state gas and to understand the CGM angular momentum and gas accretion onto galaxies.
@@
We thank the referee for thoughtful comments and suggestions that improved the manuscript. This work was supported by NASA through HST-GO-15866.002-A from the Space Telescope Science Institute. We gratefully acknowledge the support for SHH from the National Science Foundation under AST-2202328. JS was supported by the Israel Science Foundation (grant No.Ā 2584/21). Some of the data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/doi:10.17909/2b4f-h722 (catalog doi:10.17909/2b4f-h722). Some of the ground-based data collections were supported by the National Science Foundation under AST-1817125. Some observations reported here were obtained at the MMT Observatory, a joint facility of the University of Arizona and the Smithsonian Institution. Some of the data presented herein were obtained at the W.Ā M.Ā Keck Observatory, which is operated as a scientific partnership among the California Institute of Technology, the University of California and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. Some observations were supported by Swinburne Keck programmes 2014B W018E, 2015 W187E, and 2016A W032E, W056E. Some presented data are based on observations obtained with the Apache Point Observatory 3.5 m telescope, which is owned and operated by the Astrophysical Research Consortium.
@@ Appendix A Absorption System Associated with Individual Galaxies
This Appendix includes the Voigt profile fits of absorption associated with the target galaxies in individual sightlines. FigureĀ 13 (including the complete figure set in the online journal) shows the Voigt profile fits, and TableĀ 6 lists the parameters of the fits. Because not all ions were used in the analysis of this paper, we only present the measurements in the Appendix.